Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by checkley75(3666)

3C+G=1.75, G=1.75-3C
2C+4G=2.00 NOW REPLACE G WITH (1.75-3C) IN THIS EQUATION & SOLVE FOR C.
2C+4(1.75-3C)=2.00
2C+7.00-12C=2.00
2C-12C=2.00-7.00
-10C=-5.00
C=-5.00/-10.00
C=$.50 COST OF THE CANDY BAR
3*.50+G=1.75
1.50+G=1.75
G=1.75-1.50
G=$.25 COST OF GUM
PROOF
2*.50+4*.25=2.00
1.00+1.00=2.00
2.00=2.00